Use Laplace Transform to solve the following Differential Equations

a) y - 2 sin(5t) = y, y(0) = 0
Answer #1

a) 3- 2 sin(5t) - y & 4(o) = 0 Pi LCy- a sio (5t) ] = L(y) co SY (S) - YO) - 2/5 Y(S) s²+5² SY(S)- 10 = Y(s) g2+25 SY(5)-Y